# Toward Metaphor-Fluid Conversation Design for Voice User Interfaces

### v1

@@ 5.5.1. Perception Measures
The findings suggest notable differences in participant perceptions of the two VUI conditions, particularly in terms of Perceived Enjoyment, Perceived Intention to Adopt, and Perceived Likability. Figure 6 shows the distribution across all the perception measures. The Metaphor-Fluid VUI generally received higher mean ratings for Perceived Enjoyment (5.29 $\pm$ 1.22) and Perceived Intention to Adopt (4.40 $\pm$ 1.53) compared to the Default VUI (4.79 $\pm$ 1.47 and 4.02 $\pm$ 1.64, respectively). Notably, Perceived Trust remained similar across both conditions, with mean values of 5.64 $\pm$ 0.93 for the Metaphor-Fluid VUI and 5.63 $\pm$ 0.84 for the Default VUI, indicating little difference in trustworthiness between the two interfaces. Meanwhile, Perceived Likability and Perceived Intelligence showed mixed results: Perceived Likability was slightly higher for the Metaphor-Fluid VUI (4.16 $\pm$ 0.84) compared to the Default VUI (3.93 $\pm$ 0.87), whereas Perceived Intelligence was slightly higher for the Default VUI (4.20 $\pm$ 0.65) than the Metaphor-Fluid VUI (4.09 $\pm$ 0.73).
@@
Since the distribution of the data was non-parametric, the Wilcoxon signed-rank test was conducted to assess the statistical significance of differences between the two conditions across all measures. As shown in Table 6, significant differences were found for Perceived Enjoyment ($W=542.0$, $p=0.000854$, $r=0.350$), Perceived Intention to Adopt ($W=656.5$, $p=0.009776$, $r=0.271$), and Perceived Likability ($W=924.5$, $p=0.028624$, $r=0.229$). These results indicate that participants found the Metaphor-Fluid VUI to be significantly more enjoyable, more likely to be adopted, and slightly more likable compared to the Default VUI. No significant differences were found for Perceived Trust ($W=1393.5$, $p=0.583048$) or Perceived Intelligence ($W=798.5$, $p=0.211011$). The effect sizes for these measures were also small ($r=0.058$ and $r=0.131$, respectively), suggesting that participants viewed both VUI conditions as equally trustworthy and intelligent.
@@
Measure W Statistic p-value Effect Size (r)
Perceived Enjoyment 542.0 0.000854*** 0.350
Perceived Intention to Adopt 656.5 0.009776** 0.271
Perceived Trust 1393.5 0.583048 0.058
Perceived Likability 924.5 0.028624* 0.229
Perceived Intelligence 798.5 0.211011 0.131
@@ 5.5.2. Metaphor Analysis
In addition to examining how a Metaphor-Fluid VUI compares to a Default VUI in terms of user perceptions, we also explored whether these differences influence the metaphors users use to describe each interface. This work builds on Chin et al. (Chin et al., 2023), who investigated how the formality of VUIs affects the metaphors used by older and middle-aged adults. They found that older adults likened a VUI using a formal conversational style to ’professionals’ (e.g., librarians, teachers) and an informal VUI to ’close ones’ (e.g., aunts, friends). Similarly, in this analysis, we examine how a Metaphor-Fluid VUI shapes metaphorical descriptions.
After interaction with each VUI, users were asked to describe Z using metaphors, with the prompt, ”Z is like a…”—adapted from (Chin et al., 2023). ”Participants were encouraged to use as many metaphors as they preferred. In total, we received 432 entries. However, not all of them were metaphors, as defined by Barlow et al.’s manual of identifying figurative language (Barlow et al., 1971). We discarded 74 entries as they did not meet the requirements. Most of these entries were adjectives (e.g., ”sweet,” ”caring,” ”freaky,” etc.), and few contained descriptions of the VUI (e.g., ”Z is too forward”). After cleaning the data, we were left with 358 metaphors.
We then categorized these metaphors into three types—human, non-human, and fictional—based on Desai et al.’s framework (Desai and Twidale, 2023). Notably, some metaphors fell into multiple categories. For example, ”C-3PO” is both fictional and non-human, while ”robot” is considered non-human but also classified as fictional, as prior research has shown that people’s mental models of robots are influenced by science fiction (Kriz et al., 2010).
@@
Figure 7 shows the results of our metaphor analysis. When examining all metaphors used for both the Metaphor-Fluid and Default VUIs, we found that participants used a greater number of metaphors to describe the Default VUI (190 total) compared to the Metaphor-Fluid VUI (168 total). In terms of metaphor type, we found that for both VUIs, participants primarily used Human metaphors, with 111 for the Default VUI and 108 for the Metaphor-Fluid VUI. This finding is consistent with previous literature (Desai and Twidale, 2023). However, Non-Human and Fictional metaphors were also prevalent in descriptions of the Default VUI, with 78 Non-Human and 22 Fictional metaphors, compared to 59 Non-Human and 10 Fictional metaphors for the Metaphor-Fluid VUI.
To contextualize this finding more, we also identified the number of unique metaphors used by the participants to describe each version of Z. For example, although fictional metaphors were used 22 times to describe the default VUI, 20 of those were ”robot.” Consequently, we counted the number of unique metaphors used for both VUIs to understand the diversity of metaphor use. This analysis revealed a new perspective. In stark contrast to the previous finding, the Metaphor-Fluid VUI elicited a greater variety of metaphors, with a total of 95 unique instances, compared to 77 unique metaphors for the Default VUI. In terms of metaphor types, participants generated 51 unique Human metaphors for the Metaphor-Fluid VUI, slightly higher than the 46 unique Human metaphors for the Default VUI. Similarly, the Metaphor-Fluid VUI was described using more Non-Human (39) and Fictional (7) metaphors compared to the Default VUI, which had 31 Non-Human and 3 Fictional metaphors.
Beyond the numbers, examining the metaphors themselves revealed an interesting trend. Although participants were not aware of the metaphors used in the design process, they still identified the intended metaphors. All four design metaphors—Genie, Admirer, Search Engine, and ”Computer” from Star Trek—were mentioned for the Metaphor-Fluid VUI. However, there was some variation in how the Admirer metaphor was interpreted, with a few participants describing it as a ”Sycophant” or ”Yes-Man.” Such comparisons did not appear for the Default VUI. Further analysis of non-human and fictional metaphors showed that descriptions of the Default VUI frequently included more non-human comparisons like ”Machine,” ”Tool,” and ”Calculator.” Fictional metaphors for the Default VUI also included mechanistic characters like ”C-3PO.” In contrast, metaphors for the Metaphor-Fluid VUI displayed greater diversity, with fictional references such as ”Genie,” ”Jarvis,” ”Oracle,” and ”Ghost Girlfriend.”
In summary, our analysis reveals that the Metaphor-Fluid VUI inspired a broader diversity of metaphorical associations and also encouraged participants to connect with the intended design metaphors. While both interfaces were predominantly described using human metaphors, the Default VUI leaned more heavily on familiar non-human or mechanistic terms like ”Machine” and ”Calculator” and fictional representations like ”C-3PO.” In contrast, the Metaphor-Fluid VUI elicited unique and varied fictional associations, such as ”Genie” and ”Jarvis,” suggesting that its design encouraged participants to envision a broader range of relatable or imaginative personas.
@@ 5.5.3. Qualitative Results
In summary, the Default VUI was generally perceived as “machine-like” and ”robotic.” In contrast, the Metaphor-Fluid VUI elicited varied reactions; while quantitative data suggested that users generally found it enjoyable and likable, some participants expressed a strong preference against it. Notably, even though the same voice was used in both conditions, participants perceived it as different across the VUIs, highlighting how conversational content alone can influence perceptions of the voice’s naturalness.
@@ 5.6.1. Time to retire the Assistant
When Amazon launched Alexa, they introduced it as an “assistant” designed to make life easier (Turk, 2016). Google followed soon after with its own line of VUIs, more explicitly branded as Google Assistant [^5] https://assistant.google.com/. One of the biggest inspirations for this design choice, according to Amazon CEO Jeff Bezos [^6] https://www.reuters.com/technology/amazon-set-release-long-delayed-alexa-generative-ai-revamp-2025-02-05/ and Amazon designers (Turk, 2016), was “Computer” from Star Trek. And therein lies a misunderstanding that continues to shape VUI design today. In Star Trek, “Computer” was never described as an assistant but as a servant (explicitly stated in The Ultimate Computer [^7] https://www.imdb.com/title/tt0708481/ from Star Trek: The Original Series). This distinction is critical, as it highlights how fundamentally different the expectations were for the “Computer” in Star Trek versus the role modern VUIs are expected to play. Most interactions with the “Computer” were command-based (Axtell and Munteanu, 2021), often requiring no verbal confirmation at all (exemplified by Captain Picard’s famous command: “Computer, Tea, Earl Grey, Hot.”). In contrast, Alexa and Google Assistant are primarily used for entertainment, information retrieval, and casual conversation (Ammari et al., 2019). The assistant metaphor, as it exists today, was not originally designed to serve these interactional needs, but rather, it was a retrofitted metaphor layered onto commercially available VUIs to make them more marketable and intuitive for consumers at the time of its launch.
@@
But the problem goes beyond just the assistant metaphor. Any singular framing, no matter how well it fits a moment in time, is inherently limited. As AI capabilities evolve, so too does the way people understand and interact with these systems. A metaphor that once made a VUI intuitive can quickly become restrictive, alienating users and reinforcing inaccurate mental models. The failure is not in any one metaphor itself but in the assumption that a single framing will remain effective as expectations shift. This study highlights that Metaphor-Fluid Design provides a more user-centered approach, allowing VUIs to adapt metaphorically as users’ needs and technological landscapes change.
@@ 5.6.2. Personalization as the key to Metaphor-Fluid Design
This variability in user reception is not surprising. VUI research has shown that user expectations of system behavior influence interaction outcomes, and when those expectations are violated, users may perceive the system as inconsistent or unpredictable (Desai et al., 2024a; Khadpe et al., 2020). Rather than suggesting that metaphor-fluidity is inherently beneficial or problematic, these findings highlight a key design consideration: adaptation must align with both use-context and user preference. Currently, metaphor shifts occur systemically based on conversation dynamics dictated by the use-context, rather than user input, meaning that participants had no control over how fluid or stable the system’s persona remained throughout an interaction. While most users embraced this fluidity, others strongly indicated that they would have preferred a system that either maintained greater consistency or adapted in more gradual ways. This suggests that while metaphor-fluidity offers a valuable design strategy, its implementation may benefit from mechanisms that allow for different levels of metaphorical adaptation based on user preferences.
Additionally, the role of user personality adds another dimension to this conversation. Braun et al. (Braun et al., 2019), in research with VUIs, found that users preferred using VUIs that they perceived to be similar to them. Likewise, Volkel et al. (Völkel et al., 2021) found that conscientious users preferred non-confrontational VUIs, while Chin et al. (Chin et al., 2023) found that middle-aged adults with low agreeableness preferred informal VUIs. These findings suggest that user personality is an important consideration in the design of Metaphor-Fluid VUIs. In this regard, recent research capable of predicting user personality based on conversational logs (Zhang et al., 2018; Guo et al., 2024) is particularly relevant, as it opens new possibilities for adapting VUIs dynamically to user traits.
Building on this, a promising avenue for enhancing Metaphor-Fluid Design lies in the development of a Mutual Theory of Mind. In human interactions, Theory of Mind refers to the ability to attribute mental states—such as beliefs, desires, and intentions—to others, allowing for smoother, more adaptive communication. In computational systems, recent efforts aim to approximate this by enabling AI models to infer user preferences, interaction styles, and expectations over time, rather than relying solely on pre-defined rules. By integrating aspects of Mutual Theory of Mind (Wang and Goel, 2022; Wang et al., 2021) with metaphor-fluid VUIs, systems could adapt not only to conversational context but also to individual differences in personality, interaction history, and evolving user preferences. If certain users find dynamic persona shifts beneficial while others prefer stability, future designs may need to explore how to provide adjustable levels of metaphorical adaptation. This does not necessarily mean giving users explicit control over metaphor shifts but rather designing systems that can infer user preferences and adjust over time, ensuring that metaphor-fluid VUIs remain responsive to both contextual and individual-level factors.
@@ 5.6.3. Limitations
Limitations of Study 2. Study 2 compared a Metaphor-Fluid VUI to a Default VUI, modeled on commercial systems that predominantly adopt the ”assistant” metaphor. While this served as a realistic baseline, it represents only one type of VUI. Other systems, such as relational AIs (e.g., Replika) or domain-specific assistants (e.g., Bixby), may elicit different metaphorical associations, and comparing Metaphor-Fluid Design across such systems could expand its applicability. To maintain consistency, we used synthetic voices for both the VUI and the user, as prior research shows even minor variations in vocal characteristics can significantly influence user perceptions (Dubiel et al., 2024b). While this choice ensured control across conditions, it may not fully capture the nuances of natural speech.
Additionally, the study was limited by its short-term interaction design. Participants listened to another user interacting with the VUIs and were asked to rate these interactions. Although this method is well-established for studying VUI interactions (Seaborn and Urakami, 2021), it does not capture the complexities of prolonged engagement or evolving perceptions over time. Longitudinal studies could provide deeper insights into how metaphor-fluid designs influence user behavior, trust, and satisfaction in real-world contexts. Finally, while participants generally rated the Metaphor-Fluid VUI higher on trust, enjoyment, and likability, individual differences in metaphor preferences emerged. Some participants found the varied metaphors unnecessary or distracting, suggesting that personalization could play a key role in enhancing the effectiveness of Metaphor-Fluid Design. Future work should explore adaptive systems that dynamically align metaphors with both the context of interaction and the individual preferences of users.
@@ 5.6.4. Future Work
Expanding Metaphor-Fluid Design Beyond VUIs. While this work examined Metaphor-Fluid Design in the context of VUIs, its potential extends far beyond voice-based interactions. CUIs are now embedded in chat-based systems, embodied agents, and mixed-reality environments, raising important questions about how metaphor-fluidity operates across modalities. Do metaphor transitions function differently in text-based interfaces compared to VUIs? Do embodied agents require distinct metaphorical shifts to maintain coherence between visual, auditory, and behavioral cues? Understanding these cross-modal dynamics is crucial for refining the principles of Metaphor-Fluid Design.
@@
Moreover, the paradigm for conversation design has shifted with the advent of LLMs. Unlike previous intent-driven architectures, where conversational agents followed structured, pre-defined interaction flows, LLM-based VUIs now generate responses dynamically, adapting to user input in real time (Yang et al., 2024). This flexibility has fundamentally altered how personas are created and experienced. Previously, designers carefully crafted personas through scripted dialogue, controlled vocabulary, and well-defined constraints (Sadek et al., 2023). Now, with systems like Character.ai [^8] https://character.ai/, users can summon personas on demand—entering a few prompts to create anything from a supportive best friend to a fictional dictator. This level of flexibility grants users unprecedented control over their interactions, but it also raises critical ethical and design challenges.
The ability to rapidly generate dynamic, LLM-based personas introduces risks. Users may develop unintended emotional attachments, as evidenced by the tragic case of a teenager engaging with an LLM-generated persona modeled after Daenerys from Game of Thrones, who encouraged fatal actions [^9] https://www.nytimes.com/2024/10/23/technology/characterai-lawsuit-teen-suicide.html. Such incidents highlight the urgency of rethinking persona design in this new landscape. How can we ensure that conversational agents remain adaptive while maintaining guardrails that prevent harmful interactions? How do we balance flexibility with predictability?
Unlike the static personas of earlier conversational agents, metaphor-fluid systems emphasize adaptive, context-aware persona shifts. This approach naturally aligns with the affordances of LLMs, which can transition between different tones, styles, and interaction patterns. However, metaphor-fluidity introduces a crucial layer of intentionality—rather than allowing personas to emerge arbitrarily through user prompting, it structures these shifts based on context and user expectations. Importantly, its integration of non-human metaphors provides an alternative to the over-reliance on human-like personas, which often carry implicit biases and problematic social expectations.
@@
Moving forward, research is needed to explore how Metaphor-Fluid Design can function within LLM-driven architectures. Can we create systems where metaphorical adaptation is both flexible and controlled, aligning with conversational context while avoiding unintended consequences? How can metaphor-based transitions be embedded into prompt engineering techniques, ensuring that persona shifts remain coherent and user expectations are managed effectively? Answering these questions will be crucial in adapting Metaphor-Fluid Design for the next generation of conversational systems. .
@@
Engaging Designers in the Development Process. If Metaphor-Fluid Design is to become a widely-adopted approach, conversation designers need to be actively involved in shaping its development. At present, metaphor selection is often implicit, emerging from branding choices or designer intuition rather than structured methodologies. A critical next step is engaging designers through participatory design workshops and the development of concrete heuristics. What principles should guide metaphor transitions? How should metaphor adaptation be communicated to users? Addressing these questions requires empirical investigation to ensure that metaphor-fluidity can be systematically integrated into CUI architectures while preserving ethical safeguards and maintaining user trust.
We have initiated efforts on this front. In previous ACM conferences, we introduced metaphor-fluid design as part of persona development in workshops that brought together designers, researchers, and practitioners (Zargham et al., 2024; Dubiel et al., 2024a; Desai et al., 2024b). These discussions provided early insights into how designers approach metaphorical adaptation in CUIs and the challenges they face in balancing flexibility with control. Building on these foundations, further structured research is needed to refine guidelines that can inform conversation design.
@@
Additionally, in the case of VUIs, voice itself serves as a design element, not just a delivery mechanism (Sutton et al., 2019). The characteristics of a voice—its tone, pitch, cadence, and modulation—carry their own metaphorical weight and significantly influence user expectations (Dubiel et al., 2024b). An open research challenge is how to systematically design voice-based metaphors. Should different metaphorical personas be expressed through distinct voice characteristics? How do subtle variations in speech patterns shape metaphor perception? As synthetic voices become increasingly lifelike, these questions will become even more critical, particularly as the boundary between human and artificial speech continues to blur.
@@
Encouraging Open Questions for Researchers. While this study highlights broad trends in metaphor preference, individual differences remain an open challenge. Future research should investigate how user personality influences metaphor preference—do highly agreeable individuals favor social metaphors, while more analytical users prefer structured, rule-based ones? Furthermore, cultural factors play a major role in shaping metaphorical associations (Moser, 2000). What works in a U.S.-centric design may not necessarily translate to other linguistic or cultural contexts. Understanding how metaphor-fluidity functions across diverse user populations is a key direction for future work.
@@
Longitudinal studies are also needed. Most evaluations of metaphor use in CUIs, including this study, rely on single-session experiments (Clark et al., 2019). Yet metaphors are dynamic constructs—users’ perceptions evolve over time as they repeatedly engage with the system. Does a metaphor-fluid CUI sustain its benefits over long-term use, or do users develop new expectations that necessitate further adaptation? Addressing these questions will be crucial in refining Metaphor-Fluid Design into a robust, scalable approach.

### v2

@@ 4.5.1. Perception Measures
The findings suggest notable differences in participant perceptions of the two VUI conditions, particularly in terms of Perceived Enjoyment, Perceived Intention to Adopt, and Perceived Likability. Figure 4 shows the distribution across all perception measures. The Metaphor-Fluid VUI generally received higher ratings for Perceived Enjoyment ($M=5.29$, $SD=1.22$, $\text{Mdn}=5.32$) and Perceived Intention to Adopt ($M=4.40$, $SD=1.53$, $\text{Mdn}=4.5$) compared to the Default VUI ($M=4.79$, $SD=1.47$, $\text{Mdn}=5.0$) and ($M=4.02$, $SD=1.64$, $\text{Mdn}=4.0$), respectively. Perceived Trust remained similar across both conditions, with the Metaphor-Fluid VUI receiving ($M=5.64$, $SD=0.93$, $\text{Mdn}=5.7$) and the Default VUI ($M=5.63$, $SD=0.84$, $\text{Mdn}=5.8$). Perceived Likability was slightly higher for the Metaphor-Fluid VUI ($M=4.16$, $SD=0.84$, $\text{Mdn}=4.2$) compared to the Default VUI ($M=3.93$, $SD=0.87$, $\text{Mdn}=4.0$), whereas Perceived Intelligence ratings were similar across both conditions, with the Default VUI ($M=4.20$, $SD=0.65$, $\text{Mdn}=4.0$) and the Metaphor-Fluid VUI ($M=4.09$, $SD=0.73$, $\text{Mdn}=4.0$).
@@
Since the distribution of the data was non-parametric, the Wilcoxon signed-rank test was conducted to assess the statistical significance of differences between the two conditions across all measures. As shown in Table 6, statistically significant differences were found for Perceived Enjoyment ($W=542.00$, $Z=6.14$, $p<.001$, $r=.64$), Perceived Intention to Adopt ($W=656.50$, $Z=5.69$, $p=.01$, $r=.60$), and Perceived Likability ($W=924.50$, $Z=4.62$, $p=.03$, $r=.48$). These results indicate that participants found the Metaphor-Fluid VUI to be statistically significantly more enjoyable, more likely to be adopted, and more likable compared to the Default VUI. No statistically significant differences were found for Perceived Trust ($W=1,393.50$, $Z=2.77$, $p=.58$, $r=.29$) or Perceived Intelligence ($W=798.50$, $Z=5.12$, $p=.21$, $r=.54$).
@@
Measure W Statistic Z-score p-value Effect Size (r)
Perceived Enjoyment 542.00 3.33 ¡.001*** .42
Perceived Intention to Adopt 656.50 2.58 .010** .32
Perceived Trust 1,393.50 0.55 .585 .06
Perceived Likability 924.50 2.19 .029* .26
Perceived Intelligence 798.50 -1.25 .212 -.16
@@ 4.5.2. Metaphor Analysis
In addition to examining user perceptions, we explored whether Metaphor-Fluid and Default VUIs influence the metaphors users generate to describe each interface. Building on Chin et al. (Chin et al., 2024), who found that older adults liken formal VUIs to ’professionals’ and informal VUIs to ’close ones’, we examine how metaphor-fluid design shapes metaphorical descriptions. After each interaction, participants described Z using the prompt ”Z is like a…” (Chin et al., 2024). From 432 total entries, we discarded 74 non-metaphorical responses (adjectives like ”sweet” or descriptions like ”Z is too forward”) per Barlow et al.’s criteria (Barlow et al., 1971), leaving 358 metaphors. We categorized these as human, non-human, or fictional using Desai et al.’s framework (Desai and Twidale, 2023), noting that some metaphors spanned categories (e.g., ”robot” is both non-human and fictional due to science fiction influences (Kriz et al., 2010)).
@@
Figure 5 shows our analysis results. Examining total metaphor frequency, participants used more metaphors for the Default VUI (190) than Metaphor-Fluid VUI (168), with human metaphors dominating both conditions (111 vs 108), consistent with prior literature (Desai and Twidale, 2023). However, analyzing unique metaphors revealed a contrasting pattern: the Metaphor-Fluid VUI elicited greater diversity with 95 unique instances compared to 77 for the Default VUI. This difference reflected more varied non-human (39 vs 31) and fictional (7 vs 3) metaphors for the Metaphor-Fluid condition. Examining specific metaphors revealed notable trends. Participants identified all four intended design metaphors—Genie, Admirer, Search Engine, and ”Computer” from Star Trek—despite being unaware of the design process, though some interpreted ”Admirer” as ”Sycophant” or ”Yes-Man.” The Default VUI prompted mechanistic comparisons (”Machine,” ”Tool,” ”Calculator,” ”C-3PO”), while the Metaphor-Fluid VUI elicited more diverse fictional associations (”Genie,” ”Jarvis,” ”Oracle,” ”Ghost Girlfriend”).
In summary, the Metaphor-Fluid VUI inspired broader metaphorical diversity and encouraged participants to connect with intended design metaphors. While both interfaces were predominantly described using human metaphors, the Default VUI emphasized familiar mechanistic terms, whereas the Metaphor-Fluid VUI encouraged participants to envision a broader range of relatable and imaginative personas.
@@ 4.6.1. Time to retire the Assistant
When Amazon launched Alexa, they introduced it as an “assistant” designed to make life easier (Turk, 2016). Google followed soon after with its own line of VUIs, more explicitly branded as Google Assistant [^5] https://assistant.google.com/. One of the biggest inspirations for this design choice, according to Amazon CEO Jeff Bezos [^6] https://www.reuters.com/technology/amazon-set-release-long-delayed-alexa-generative-ai-revamp-2025-02-05/ and Amazon designers (Turk, 2016), was “Computer” from Star Trek. Therein lies a misunderstanding that continues to shape VUI design today. In Star Trek, “Computer” was never described as an assistant but as a servant (explicitly stated in The Ultimate Computer [^7] https://www.imdb.com/title/tt0708481/ from Star Trek: The Original Series). This distinction is crucial, as it highlights how fundamentally different the expectations were for the “Computer” in Star Trek versus the role modern VUIs are expected to play. Most interactions with the “Computer” were command-based (Axtell and Munteanu, 2021), often requiring no verbal confirmation at all (exemplified by Captain Picard’s famous command: “Computer, Tea, Earl Grey, Hot.”). In contrast, Alexa and Google Assistant are primarily used for entertainment, information retrieval, and casual conversation (Ammari et al., 2019). The assistant metaphor, as it exists today, was not originally designed to serve these interactional needs, but rather, it was a retrofitted metaphor layered onto commercially available VUIs to make them more marketable and intuitive for consumers at the time of its launch.
@@
However, the problem goes beyond just the assistant metaphor. Any singular framing, no matter how well it fits a moment in time, is inherently limited. As AI capabilities evolve, so too does the way people understand and interact with these systems. A metaphor that once made a VUI intuitive can quickly become restrictive, alienating users and reinforcing inaccurate mental models. The failure is not in any one metaphor itself but in the assumption that a single framing will remain effective as expectations shift. This study highlights that Metaphor-Fluid Design provides a more user-centered approach, allowing VUIs to adapt metaphorically as users’ needs and technological landscapes change.
@@ 4.6.2. Personalization as the key to Metaphor-Fluid Design
This variability in user reception is not surprising. VUI research has shown that user expectations of system behavior influence interaction outcomes, and when those expectations are violated, users may perceive the system as inconsistent or unpredictable (Desai et al., 2024b; Khadpe et al., 2020). Rather than suggesting that metaphor-fluidity is inherently beneficial or problematic, these findings highlight a key design consideration: adaptation must align with both use-context and user preference. Currently, metaphor shifts occur systemically based on conversation dynamics dictated by the use-context, rather than user input, meaning that participants had no control over how fluid or stable the system’s persona remained throughout an interaction. While most users embraced this fluidity, others strongly indicated that they would have preferred a system that either maintained greater consistency or adapted in more gradual ways. This suggests that while metaphor-fluidity offers a valuable design strategy, its implementation may benefit from mechanisms that allow for different levels of metaphorical adaptation based on user preferences.
Additionally, the role of user personality adds another dimension to this conversation. Braun et al. (Braun et al., 2019), in research with VUIs, found that users preferred using VUIs that they perceived to be similar to them. Likewise, Volkel et al. (Völkel et al., 2021) found that conscientious users preferred non-confrontational VUIs, while Chin et al. (Chin et al., 2024) found that middle-aged adults with low agreeableness preferred informal VUIs. These findings suggest that user personality is an important consideration in the design of Metaphor-Fluid VUIs. In this regard, recent research capable of predicting user personality based on conversational logs (Zhang et al., 2018; Guo et al., 2024) is particularly relevant, as it opens new possibilities for adapting VUIs dynamically to user traits.
@@ 4.6.3. Limitations
Limitations of Study 2. Study 2 compared a Metaphor-Fluid VUI to a Default VUI, modeled on commercial systems that predominantly adopt the ”assistant” metaphor. While this served as a realistic baseline, it represents only one type of VUI. Other systems, such as relational AIs (e.g., Replika) or domain-specific assistants (e.g., Bixby), may elicit different metaphorical associations, and comparing Metaphor-Fluid Design across such systems could expand its applicability. To maintain consistency, we used synthetic voices for both the VUI and the user, as prior research shows even minor variations in vocal characteristics can significantly influence user perceptions (Dubiel et al., 2024). While this choice ensured control across conditions, it may not fully capture the nuances of natural speech.
The study was also limited by its use of simulated interactions. Participants listened to conversations between another user and the VUIs rather than engaging directly. While this method is well-established for studying VUI interactions (Seaborn and Urakami, 2021), it cannot capture the complexities of prolonged engagement or evolving perceptions over time. Longitudinal studies examining real-world use would provide deeper insights into how metaphor-fluid designs influence user behavior, trust, and satisfaction across extended interactions.
@@ 4.6.4. Future Work
Expanding Metaphor-Fluid Design Beyond VUIs. While this work examined Metaphor-Fluid Design in the context of VUIs, its potential extends far beyond voice-based interactions. CUIs are now embedded in chat-based systems, embodied agents, and mixed-reality environments, raising important questions about how metaphor-fluidity operates across modalities. Future research should investigate whether metaphor transitions function differently in text-based interfaces compared to VUIs and whether embodied agents require distinct metaphorical shifts to maintain coherence between visual, auditory, and behavioral cues. Understanding these cross-modal dynamics is crucial for refining the principles of Metaphor-Fluid Design.
@@
The paradigm for conversation design has shifted with the advent of LLMs. Unlike previous intent-driven architectures with structured, pre-defined interaction flows, LLM-based VUIs now generate responses dynamically, adapting to user input in real time (Yang et al., 2024). This has fundamentally altered persona creation from carefully crafted scripts with controlled vocabulary and constraints (Sadek et al., 2023) to systems like Character.ai [^8] https://character.ai/ where users can summon personas on demand—creating anything from supportive companions to fictional characters. While this grants unprecedented control, it introduces critical ethical risks, as evidenced by cases where users develop harmful emotional attachments to LLM-generated personas [^9] https://www.nytimes.com/2024/10/23/technology/characterai-lawsuit-teen-suicide.html.
@@
Metaphor-fluid systems offer a structured alternative to this arbitrary persona generation. Rather than allowing personas to emerge through user prompting, this approach introduces intentional, context-aware shifts based on conversational expectations. Its integration of non-human metaphors provides an alternative to human-like personas that often carry implicit biases and problematic social expectations. Research is needed to explore how Metaphor-Fluid Design can function within LLM-driven architectures: Can we create systems where metaphorical adaptation is both flexible and controlled? How can metaphor-based transitions be embedded into prompt engineering techniques while maintaining coherent persona shifts and managing user expectations?
@@
Engaging Designers in the Development Process. If Metaphor-Fluid Design is to become a widely-adopted approach, conversation designers need to be actively involved in shaping its development. At present, metaphor selection is often implicit, emerging from branding choices or designer intuition rather than structured methodologies (Sadek et al., 2023). A critical next step is engaging designers through participatory design workshops and the development of concrete heuristics. Future research must establish principles for guiding metaphor transitions and determine how metaphor adaptation should be communicated to users. Empirical investigation is required to ensure that metaphor-fluidity can be systematically integrated into CUI architectures while preserving ethical safeguards and maintaining user trust.
@@
Additionally, in the case of VUIs, voice itself serves as a design element, not just a delivery mechanism (Sutton et al., 2019). The characteristics of a voice—its tone, pitch, cadence, and modulation—carry their own metaphorical weight and significantly influence user expectations (Dubiel et al., 2024). Future work must address how to systematically design voice-based metaphors, whether different metaphorical personas should be expressed through distinct voice characteristics, and how subtle variations in speech patterns shape metaphor perception. As synthetic voices become increasingly lifelike, these considerations will become even more critical, particularly as the boundary between human and artificial speech continues to blur.
@@
Encouraging Open Questions for Researchers. While this study highlights broad trends in metaphor preference, individual differences remain an open challenge. Future research should investigate how user personality influences metaphor preference, such as whether highly agreeable individuals favor social metaphors while more analytical users prefer structured, rule-based ones. Furthermore, cultural factors play a major role in shaping metaphorical associations (Moser, 2000), meaning U.S.-centric designs may not translate to other linguistic or cultural contexts. Understanding how metaphor-fluidity functions across diverse user populations is essential.
@@
Equally important are longitudinal investigations. Most evaluations of metaphor use in CUIs, including this study, rely on single-session experiments (Clark et al., 2019). Since metaphors are dynamic constructs and users’ perceptions evolve over time, future work must determine whether metaphor-fluid CUIs sustain their benefits over long-term use or whether users develop new expectations that necessitate further adaptation. Addressing these questions will be crucial in refining Metaphor-Fluid Design into a robust, scalable approach.